OpenStudy (anonymous):

Find the volume of a cylinder in terms of 'r' if its height is five times the size of its radius. Volume of a cylinder is: V = ♊*r^2*h

OpenStudy (anonymous):

h=5r v=pi*r^2*5r

OpenStudy (anonymous):

Hey Nukka, my answer book say the answer is: V = 5*pi*r^3

OpenStudy (anonymous):

Yeah i used volume of a rectangle formula.. Vcyl= pi*r^2*h

OpenStudy (anonymous):

Oh I see, thanks

OpenStudy (anonymous):

erm... no the answer was right. Just needed to expand it... Looks like its time to sleep.

OpenStudy (anonymous):

Haha thankyou very much